Calculate the electrostatic force of attraction between the sole electron and the nucleus of the hydrogen atom. [ Take electron charge = 1.6×10^-19C , Proton Charger = 1.6×10^-19C , Radius of hydrogen atom = 10^-10 m , K= 9.0×10^-9NM^2C^-2 ]
1
Expert's answer
2020-02-12T10:04:52-0500
"F=\frac{ke^2}{r^2}"

"F=\frac{(9.0\cdot10^{9})(1.6\cdot10^{-19})^2}{(10^{-10})^2}=2.3\cdot10^{-8}N"